OpenStudy (anonymous):

Can someone please help me solve this problem: solve the equation log(1-x)=log(10) for x?

OpenStudy (anonymous):

1-x=10 -x=10-1 -x=9 x= -9

OpenStudy (anonymous):

Thank you (: